The base of a triangle is one-third as long as its height. If the area of the triangle is 24 square meters, how long is its base?
Found 2 solutions by mananth, rfer:
Answer by mananth(16949)

let height be x meters
base = 1/3 x
.
area of traingle = 1/2 * base * height
= 1/2 * 1/3 *x *x
=x^2/6 sq. meters

Answer by rfer(16322)

x*3x=48
3x^2=48
x^2=16
x=4 meters
3x=12 meters